please help me in this problem.
Mark is two-thirds as old as Helen. In seven years, Mark will be three-fourths as old as Helen. How old are they now?
Thanks!

Answer by J2R2R(94)

For now, Mark is 2/3 of Helen, so let us say Mark is 2x, therefore Helen is 3x.
In seven years, 2x+7 = 3(3x+7)/4 or 4(2x+7) = 3(3x+7)
Giving 8x+28 = 9x+21
Solved gives x = 7
Let us test it:
Now: Mark is 14 and Helen is 21 (two thirds of 21 = 14)
Add 7: Mark will be 21 and Helen will be 28 (21 is three quarters of 28)